FEL Upgrade Project Weekly Brief
August 13-17, 2001




Highlights:
This week Larry Carr (Brookhaven National Lab), Michael Martin and Wayne McKinney (Lawrence Berkeley National Lab), joined G. Neil, G. Williams, M. Shinn and K. Jordan to set up a rapid-scan Michelson interferometer to observe multiparticle coherent synchrotron radiation in the THz (far-IR) spectral region.  They had a successful run.  They observed far-IR coherent synchrotron radiation whose intensity varied as the charge per bunch squared.  They measured its spectral content, and also the absolute power.  The effects of bunch length were also measured.   In this spectral region such a source provides almost 1watt/cm-1 of average power, which is many orders of magnitude higher than any other source to date and probably enough to drive non-linear processes in a new regime.

This work is important to the 3 laboratories in 2 respects.  It links to the many projects involvingaccelerator- produced IR and also sheds light on the utility of Energy Recovered Linac machines, mainly in addressing critical stability questions.

Management:
Due to management participation at the FEL Conference, August 19-24, 2001 in Darmstadt, Germany there will not be a weekly report next week.

WBS 4 (Injector):
Signed off the ball cathode drawing and submitted the job along with the material to the shop. Ordered and received the material for screens in the gun chamber.

Assembled and tweaked the cathode shield assembly (dirty), parts are being cleaned and some plated prior to assembly into a test vacuum system.

WBS 6 (RF):
Zone 3 - No Progress.  Waiting to get into the vault after the summer operations.

Zone 4 - No Progress.  Recovering missing CAMAC boards.

Gun HVPS - Reviewing proposed schedule with intent of making it shorter by 3 months.  Some tasks may have to be postponed until later.  The Mechanical Engineer and Designer will make field measurements after the summer operations.

Quarter Klystrons - CPI has started the final bake and evacuation of the first klystron.  They are intend to have factory tests witnessed on 9/12/01.

Quarter HVPS's - The 3 Circulators are scheduled to be tested at 50 kW next week after the summer operations.  The
representative from AFT, the vendor, will witness and assist in these tests.  The circulators will be tested at 100 kW after the new 100 kW klystrons and HVPS's have been installed and checked out.  The SCR Controller and Contactor were ordered this week.  Schematics and Parts Lists are being made.

WBS 8 (Instrumentation):
Testing is nearly complete on the VME timing/delay board and EPICS software is being developed to support it.   Prototype boards for the new beamviewer camera/lamp board power supplies have arrived and all components for the beamviewer control box are on order.  Additional carrier boards have been ordered to be used as new systems are developed.

An analog differential driver board is also under development for better monitoring of beam current.

The first of the new Ion Pump Power Supplies has been shipped back to the lab for testing.

The BAIR web and the task system are continually being upgraded to make them more useful.  A digital Input/Output box is being built for use in the labs to allow more functionality and control of test equipment.

Drawings completed this week:  F0015 - 32 Channel Sample and Hold Board drawings have been revised to correct the
placement of the connectors.  F0060 - VME Timing Main Board was revised to make a minor change in hole size for the connector and one of the front test connections.  F0094-98 - Temperature Monitoring Chassis:  Wiring Diagram, Fabrication, and Assembly Drawings

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
o  We are granting a one week extension to the Aug. 20 bidding period because some of our usual vendors have not
    been able to get subcontractor bids in time.
Injector Dipoles (DU/DV)
o  The first round of rework on some Assembly Drawings were received and commented on.  We also defined the
    wording and graphics for the Purcell gap gluing for application to the GU assembly drawings.
Arc Dipoles (GY, GX, GQ)
o  In magnetic modeling of the 180 Dipole (GY), AES completed their analysis, nominally achieving David Douglasí
    specifications.  They are submitting those values to Dave for final approval. (Two down, one to go)
o  They started and completed the magnetic model of the Reverse Bend (GQ), extracting many of the features of the
    First Bend (GX) into this simpler model.  They will start their computer runs to discover if it falls within magnetic
    specification over the weekend.
o  Design continued on back checking the detail drawings of all three magnets with the comments and corrections that
    have built up over the course of the project.  Work will begin next week on bringing the DY drafting model (and
    hence, automatically most of the details) up to the dimensions of the successful magnetic model.
Quadrupoles
3 inch quad (QX)
o  The steel for cores was received by New England Technicoil.
o  Al Guerra and Robin Wines inspected the first article of the quad coil at Magnet Enterprises International on
    Oakland CA.
Trim Quad (QT)
o  The 2D detail drawings have gone through one cycle of checking and backchecking and are ready for a review by
    the group. The Statement of Work was composed and is in review.
Sextupole (SF)
o  Drawing layout continued.
Octupole (OT)
o  DULY Research started their magnetic modeling.
Corrector Dipoles (DB, DJ)
o  Tom Hiatt visited Milhous Co. to assess progress in manufacturing the corrector coil.  They were found to be
    working out tensioning and potting bugs, but their manufacturing tooling and plan is in order.
o  We detailed the design of the new mounting bracket that will hold the DB/DJ correctors over the new larger
    diameter beam position detectors.
Beam Line and Vacuum
o  Work continued on the definition of the Optical Chicane chambers.
o  The bid package drawing set for the girders in the return leg continues in detailing.
o  We decided to use the recycled stands in the UV line and use new ones in the IR line to cut down the cycle time
    between tear out and installation.
o  We started engineering work on the LCW distribution again to respond to the new availability of the installation
    crew in the early FY 02 period.
o  We found there is still a controversy on the style of the final dump.  George Neil is still an advocate of the garbage
    can style catch beam even without an optimal set up while Dave Douglas is advocating a narrow design that he
    claims provides the diagnostics for a required set up and is ultimately more robust because of its predictability.
General
o  The Fast Track Schedule is now up to date and has the greater detail necessary to give a better view of how the
    project is progressing.

WBS 10 (Wiggler):
Work on the wiggler magnet test stand concentrated on device calibration.  The test magnet originally planned for use in probe calibration was found to be unacceptable so the BA magnet will be used.

Progress was made on calibrating the motion encoder on the wiggler magnet measurement apparatus.  Another day or two of tests should allow the magnet measurements to start.  The drawing of the correction coil for the dispersion section was released so that work could begin on the coil installation.

WBS 11 (Optics):
Design of the optical cavity chambers is at the point where we feel confident we can purchase some components.  Creation of a parts list is underway, and we plan to place orders next week.  The engineer assigned to assist in the design of OCMMS (optical cavity alignment and metrology) hardware has a good design and we've asked for a designer to refine it.

We've sent responses to questions from respondents to the drive laser solicitation.

Support for operations/experimenters:

On 8/10, did measurements to determine the loss of one of our 6 um HR mirrors.  Preliminary analyses of the data sets the loss below 300 ppm.  Provided support for the primary users (THz generation), and participated in the imaging experiments.  A highlight was the imaging of the shadow of a smartcard chip embedded in it's plastic holder.  Analyzed epoxy tests, determining the bond thickness to be 175 microns of the epoxy we will use to affix mirrors to the deformable mirror assembly.  Outgassing wasn't excessive; pumping overnight reduced this to below 1e-9 Torr.

Operations/Commissioning:
This week 3 different user groups performed their experiments with the machine.

Measurements of the higher-order mode power in the electron beam and its dependence on beam parameters were made on Monday & Tuesday.  A complete set of data was obtained analysis.

The THz radiation experiments were made on Wednesday and Thursday with visiting scientists from BNL and LBNL.  This successful experiment is described in the Highlights sections of this report.

Rapid thermal processing of polymers was conducted on Friday.  The earlier problems with this experiment were solved and it successfully ran.  Differences in macropulse amplitude stability at high rep. rates are being investigated.

This is the last week of operations for the Summer Run.  The machine will be opened next week.
 


Privacy and Security Notice